OpenStudy (anonymous):

Really confused! please help! The density of a fish tank is 0.5fish over feet cubed. There are 30 fish in the tank. What is the volume of the tank? 60 ft3 120 ft3 6 ft3 15 ft3

OpenStudy (fibonaccichick666):

so, p=m/v ie. density equals mass over volume. so in this case the "mass" is the number of fish. are you able to solve \[\rho=\frac{m}{v}\] for v, then plug in the values for rho and m?

OpenStudy (anonymous):

im sorry.. i moved to a new school so i havnt learned this yet..

OpenStudy (fibonaccichick666):

oh, well, the formula, you just have to memorize. As for solving, you just solve the formula for the variable v then plug and chug.

OpenStudy (fibonaccichick666):

it is just like solving for y or x when you have the equation for a line

OpenStudy (anonymous):

so would it be B?

OpenStudy (alekos):

No, write out the equation and solve for v

OpenStudy (anonymous):

i dont think im getting it right...

OpenStudy (fibonaccichick666):

Please show your work solving for v here

OpenStudy (fibonaccichick666):

it should take two steps

OpenStudy (anonymous):

the only formulas i know are formulas like midpoint and distance.. im not sure how to do what your asking.. i dont think ive learned this at all

OpenStudy (fibonaccichick666):

I told you the formula. move the variables around to solve for v

OpenStudy (anonymous):

p=30/v?

OpenStudy (fibonaccichick666):

first you need to make the formula look like v=...

OpenStudy (fibonaccichick666):

although, mass will equal 30

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

so.. is the density of the fish v?

OpenStudy (fibonaccichick666):

no, density is density (p or rho)

OpenStudy (anonymous):

ohhhh ok... sorry.. im not very good with math haha

OpenStudy (anonymous):

@Directrix sorry for asking you to help me so much..

OpenStudy (anonymous):

i think its 15 cubed.. am i right?

Directrix (directrix):

I don't think so.

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

i multiplied 0.5 by 30 and solved to get volume

OpenStudy (anonymous):

wait.. i know what i did wrong!

OpenStudy (anonymous):

i got 60..

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

yes 60 is correct.

OpenStudy (anonymous):

yay! could you help me with another question?

Directrix (directrix):

Start a new thread.

OpenStudy (anonymous):

ok :)